# CANISIUS The Austrian Neutron Spin Echo Interferometer

### v1

@@ II.5 Coherent Averaging
Constructing structured wavefunctions by coherent averaging is the main purpose of the CANISIUS instrument. This is done primarily to produce neutron wavefunctions that have non zero Orbital Angular Momentum (OAM) and investigate the properties of said states. We have demonstrated the method previously in perfect crystal neutron interferometry [10]. Here the structured wave is produced by running the instrument in a SESANS like mode, where the first and second arm have different poleshoe angles. As a result the two spin states retain some transverse separation after passing through the spin echo interferometer. The concept is illustrated in figure 8. This residual transverse separation is given by
@@
with $\sigma$ the transverse coherence length the output wavefunction after spin selection would be equal to
$\psi_{1}=A^{\prime}[e^{-(x-\frac{\delta_{res}}{2})^{2}/\sigma^{2}}+e^{i\eta}e^{-(x+\frac{\delta_{res}}{2})^{2}/\sigma^{2}}]e^{ik_{y}y}$ (16)
with $\eta$ a longitudinal phase between the two partial wavefunctions. $\eta$ can be precisely tuned by changing the phase of one of the RF coils $\eta=2\Delta\phi$. If $\delta_{res}$ is chosen appropriately with respect to the transverse coherence length $\sigma$ and $\eta=\pi$, the wavefunction could be approximated by
$\psi_{1}^{\prime}\propto\Pi(\frac{2k_{x}x}{\pi})\sin(k_{x}x)e^{ik_{y}y}$ (17)
such a wavefunction mimics a linearly polarized oscillation, which can also can be described by two counter rotating oscillations. This can easily be shown by describing the above equation in cylindrical coordinates.
$\psi_{1}^{\prime}\propto\frac{\mathrm{i}}{2}\Pi(\frac{2k_{x}x}{\pi})[e^{-ik_{x}r\cos{\phi}}-e^{ik_{x}r\cos(\phi)}]e^{ik_{y}y}$ (18)
The exponential functions that have replaced the sine function can be expanded according to the Jacobi Anger expansion [46].
$\psi_{1}^{\prime}\propto\frac{\mathrm{i}}{2}\Pi(\frac{2k_{x}x}{\pi})[\sum_{\ell}J_{\ell}(-k_{x}r)e^{\mathrm{i}\ell\phi}-J_{\ell}(k_{x}r)e^{\mathrm{i}\ell\phi}]e^{ik_{y}y}$ (19)
Since Bessel functions are even for the zeroth mode and all even mode numbers $\ell$ this reduces to
$\psi_{1}^{\prime}\propto-\mathrm{i}\Pi(\frac{2k_{x}x}{\pi})\sum_{\ell=\mathrm{odd}}J_{\ell}(k_{x}r)e^{\mathrm{i}\ell\phi}e^{ik_{y}y}$ (20)
The rectangle function ensures that the $\psi_{1}^{\prime}$ is only non zero for small arguments of the Bessel function. Since the Bessel function of order $\ell$ is proportional to $z^{ell}$ (with z the argument), only the first order modes are relevant.
$\psi_{1}^{\prime}\propto-\mathrm{i}\Pi(\frac{2k_{x}x}{\pi})\sum_{\ell=\pm 1}J_{\ell}(k_{x}r)e^{\mathrm{i}\ell\phi}e^{ik_{y}y}$ (21)
Similarly it can be shown that for a phase $\eta=0$ the $\ell=0$ and =±$\ell=\pm 2$ modes are generated. Even though the average OAM of each wavefunction produced by CANISIUS by coherent averaging is zero, the superpositions of =±$\ell=\pm 1$ or ±$\pm 2$ are qualitatively different than the $\ell=0$ mode, which are expected to scatter differently from nuclear targets [12, 13]. In addition the absorption characteristiscs of these states are expected to be different [14]. And as described in a recent paper these states may be used to explore the quantum Sagnac effect [47].
@@ III Conclusion
We have presented the new Austrian spin echo interferometer of the Atominstitut, CANISIUS and illustrated its capabilities, both as a ToF and a broad band white beam instrument. In particular we demonstrated the feasibility of polychromatic SESANS, which allows us to use the full potential of a continuous neutron source. This is especially important at weaker sources such as the 250 kW Triga reactor at the Atominstitut. Finally the coherent averaging capabilities of the instrument were illustrated. It was shown that CANISIUS is capable of producing neutrons that carry non zero OAM, in particular states that exist in a superpostion of the =±$\ell=\pm 1$ modes. We expect this will be useful for exploring OAM dependent scattering and absorption cross sections, in addition to exploring fundamental aspects of quantum mechanics from the quantum Sagnac effect to quantum information and contextuality.
@@ IV Acknowledgments
N. G. and S. S. acknowledge funding from the Austrian science fund (FWF), Project No. P34239, in addition N. G. is supported by the US Department of Energy (DOE) grant DE-SC0023695.

### v2

@@ II.5 Coherent Averaging
Constructing structured wavefunctions by coherent averaging is the main purpose of the CANISIUS instrument. This is done primarily to produce neutron wavefunctions that have non zero Orbital Angular Momentum (OAM) and investigate the properties of said states. Put simply coherent averaging is an interferometric technique that takes an input wavefunction with simple structure (i.e. Gaussian) and splits it multiple identical copies, called partial wavefunctions. These partial wavefunctions can then be independently translated and phase shifted in real or reciprocal space to produce a composite wavefunction (i.e. the coherent sum of all partial wavefunctions) that exhibits the desired structure. This technique is particularly well suited for producing neutron OAM, as explored in [48, 9, 11, 10]. Coherent averaging could be used to produce a circular array of partial wavefunctions which act as coherent sources. If one can control the phase of each source, it becomes possible to produce arbitrary OAM values as described in [49]. We have demonstrated the method previously in perfect crystal neutron interferometry [10].
Since SESANS is also a type of neutron interferometer it can be used to produce structured wavefunctions by coherent averaging. Here the structured wave is produced by running the instrument in a new mode, where the first and second arm have different poleshoe angles. As a result the two spin states retain some transverse separation after passing through the spin echo interferometer. The concept is illustrated in figure 8. This residual transverse separation is given by
@@
with $\sigma$ the transverse coherence length the transvers part of the output wavefunction after spin selection would be equal to
$\psi_{1}=\frac{A}{\sqrt{2}}(e^{\mathrm{i}\beta/2}e^{-\frac{(x-\delta_{res}/2)^{2}-y^{2}}{\sigma^{2}}}+e^{-\mathrm{i}\beta/2}e^{-\frac{(x+\delta_{res}/2)^{2}-y^{2}}{\sigma^{2}}})$ (16)
with $\beta$ a longitudinal phase between the two partial wavefunctions. $\beta$ can be precisely tuned by changing the phase of one of the RF coils $\beta=2\Delta\phi$. We can analyze the OAM or mode distribution by introducing the azimuthal Fourier transform
$\psi^{\ell}(r,z)=\frac{1}{2\pi}\int_{0}^{2\pi}\psi(r,\phi,z)e^{-\mathrm{i}\ell\phi}\mathrm{d}\phi$ (17)
and its inverse
$\psi(r,\phi,z)=\sum_{\ell}\psi^{\ell}(r,z)e^{\mathrm{i}\ell\phi}$ (18)
When we apply the azimuthal Fourier transform to our composite wavefunction we find
$\psi_{1}^{\ell}(r)=\sqrt{2}Ae^{-\frac{r^{2}+\delta_{res}^{2}/4}{\sigma^{2}}}J_{\ell}(\mathrm{i}\frac{\delta_{res}}{\sigma^{2}}r)\cos(\frac{\pi\ell+\beta}{2})$ (19)
Since this mode amplitude is complex and spatially dependent it can be useful to introduce a mode distribution function defined as
$p[\ell]=2\pi\int r|\psi^{\ell}(r)|^{2}\mathrm{d}r$ (20)
which is related to the OAM expectation value through
$\hat{L}_{z}>=\hbar\sum_{\ell}\ell p[\ell]$ (21)
When we calculate the mode distribution function of 19 we find
$p[\ell]=\frac{A^{2}\sigma^{2}}{2}e^{-\frac{\delta_{res}^{2}}{4\sigma^{2}}}I_{\ell}(\frac{\delta_{res}^{2}}{4\sigma^{2}})|\cos|^{2}(\frac{\pi\ell+\beta}{2})$ (22)
We may now observe that for $\beta=0$, $p[\ell]$ is only non zero for even $\ell$, that is to say only the even OAM modes contribute to the output wavefunction, while for $\beta=\pi$ $p[\ell]$ is only non zero for odd $\ell$ and hence only odd OAM modes contribut to the output wavefunction. By tuning the ratio $\delta_{res}/\sigma$ we can determine how many modes contribute. For small splitting only few modes contribute whereas for large splitting many modes shape the wavefunction. The OAM distribution function along with these properties are shown in figure 9. Interestingly we find that for $\beta=\pi$ and small $\delta_{res}$ with respect to the coherence length the output wavefunction can almost entirely be described by the $\ell=-1$ and $\ell=1$ modes, i.e. a right and left twisting mode, somewhat analogous to linear polarization. The concept of a ”linear” OAM state is illustrated in figure 10.
Figure 10: Illustration of the construction of a linearly polarized OAM state. A clockwise rotating state (a)/(d) has a counter clockwise rotating state (b)/(e) subtracted from it, producing a linearly polarized state (c)/(f). The top insets (a)-(c), show the imaginary parts of the respective transverse wavefunction, while the bottom insets (d)-(f) show the absolute values, which is directly related to the expected intensity pattern.
Here a left rotating state is subtracted from a right rotating state to produce a linearly polarized (sine oscillation) state. We have now shown that such a linearly polarized state can be approximated by a superposition of two Gaussian wavepackets with opposite amplitudes on opposite sides of the cylinder axis (eq. 16). If $\delta_{res}$ is smaller than or on the order of $\sigma$, we find that figure 9 and equation 22 demonstrate that this very closely approximates a linearly polarized OAM state with $|\ell|=1$. While this state has an average OAM of zero just like a pure $\ell=0$ mode it is orthogonal to a pure $\ell=0$ state as nicely illustrated by looking at the OAM distribution function. In addition, even though the average OAM of each wavefunction produced by CANISIUS by coherent averaging is zero, the superpositions of $\ell=\pm 1$ or $\pm 2$ are qualitatively different than the pure $\ell=0$ mode, which are expected to scatter differently from nuclear targets [12, 13]. In addition the absorption characteristics of these states are expected to be different [14]. And as described in a recent paper these states may be used to explore the quantum Sagnac effect [50], where these superposition states may produce intensity beating proportional to the mode number and the rotation frequency of the observer.
@@ III Conclusion
We have presented the new Austrian spin echo interferometer of the Atominstitut, CANISIUS, and illustrated its capabilities, both as a ToF and a broad-band white beam instrument. In particular we demonstrated the feasibility of polychromatic SESANS, which allows us to use the full potential of a continuous neutron source. This is especially important at weaker sources such as the 250 kW Triga reactor at the Atominstitut. Finally the coherent averaging capabilities of the instrument were illustrated. It was shown that CANISIUS is capable of producing neutrons that carry non zero OAM, in particular states that exist in a superpostion of the $\ell=\pm 1$ modes. We expect this will be useful for exploring OAM dependent scattering and absorption cross sections, in addition to exploring fundamental aspects of quantum mechanics from the quantum Sagnac effect to quantum information and contextuality.
@@ IV Acknowledgments
N. G. and S. S. acknowledge funding from the Austrian science fund (FWF), Project No. P34239, in addition N. G. is supported by the US Department of Energy (DOE) grant DE-SC0023695. The thermal white beam facility was designed, constructed and funded by the Federal Ministry of Education, Science and Research of Austria. Finally the authors acknowledge and thank Wim Bouwman for useful discussion and resources regarding white beam SESANS.